OpenStudy (anonymous):

Heart failures are due to either natural occurrences (87%) or outside factors (13%). Outside factors are related to induced substances (73%) or foreign objects (27%). Natural occurrences are caused by arterial blockage (56%), disease (27%), or infection (17%). a) Determine the probability that a failure is due to induced substance. b) Determine probability that failure is due to disease or infection.

OpenStudy (anonymous):

you, a)Lets denote probability due induce p(I)=0.73x0.13 b) Probability due to diseas or infection = P(d)Up(i)=P(d)+p(i)=(0.17+0.27)x0.87